Naturally occurring cobalt consists of only one isotope 59 Cowhose relative atomic mass is 58.9332. A synthetic radioactive isotope of cobalt 60 Co has relative atomic mass 59.9338, is used in radiation therapy for cancer. A 1.8516 sample of cobalt has an apparent "atomic mass" of 58.9901. Find the mass of 60 Co is this sample?
1
Expert's answer
2012-07-13T09:35:25-0400
58.9332*x + 59.9338*(1 – x) = 58.9901
59.9338 – 1.0006*x = 58.9901
1.0006*x = 0.9437
x = 0,94313 = 94.313%
1 – x = 0.05687
The mass of 60 Co will be:
1.8516*0.05687 = 0.1053 g